OpenStudy (anonymous):

-2+3(7V+6)=6(6v-8)+4

OpenStudy (anonymous):

hold on doing the work now

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

is there multiple choice?

OpenStudy (anonymous):

no, its just the equation on my worksheet

OpenStudy (anonymous):

i just need someone to kinda walk me through it is all, like whats the first step?

OpenStudy (anonymous):

-2+21v+18=36v-48+4 16+21v=36v-44 60+21v=36v 60=15v 4=v

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

Anything else?

OpenStudy (anonymous):

nah im good, thanks

OpenStudy (anonymous):

wait come back

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

im not getting 16+21v=36v-44 i keept getting: 21v+20=36v-44

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

kidding, i fixed it.

OpenStudy (anonymous):

well i was almost done but okay bye